# Incorporating Magnetic Field Characteristics into EUV-Based Automated Segmentation of Coronal Holes

### v1

@@ 4.2 Determining Region Evolution
In order to determine if filament regions can be removed during evolution, instead of as a post-hoc process, the unipolarity of every region as a function of number of iterations was calculated for both the EUV-only and EUV + HMI magnetogram segmentations. This was achieved by saving a copy of the contour after each iteration (assuming seeding and iteration at one-eighth-scale resolution). Once the output of each iteration was obtained, the segmentations were resized to the full-scale resolution, and the line-of-sight unipolarity of each region was calculated using the original, full-scale HMI magnetogram observation.
Figure 13 provides two examples of the change in measured unipolarity of a filament region as a function of ACWE evolution and the corresponding change in area measured in pixels at the full-scale resolution. For both examples we provide the line-of-sight unipolarity, calculated from the HMI magnetogram via Equation 6, and a weighted unipolarity, weighting regions near disk center (where estimates of magnetic field strength are more accurate) using the method outlined in the SunPy documentation (The SunPy Community, 2024) as functions of the number of iterations on the left panel of each example. In the right panel of each example, we provide the area of the region as a function of the number of iterations.
The two examples in Figure 13 were chosen to demonstrate the two primary behaviors noted among filaments in CR 2133. The first case, seen in Figure 13a, consists of regions that quickly become bipolar as they evolve. In these cases, minimal evolution, even with spatially-decimated HMI magnetogram data, causes the region to quickly become bipolar. We note that when this occurs, the spatially-decimated HMI magnetogram further constrains evolution as compared to an EUV-only evolution, preventing further growth of the region. In these cases the filament region can be identified and removed after very few iterations. This case is representative of most filaments found in CR 2133.
@@
Figure 13: Region unipolarity (left) and region size in pixels (right) as a function of iteration for EUV-only and EUV+HMI magnetogram segmentations generated at one-eighth scale resolution.
The second case, represented in Figure 13b, consists of regions that take several iterations before they can be correctly identified as filament regions. In these cases the spatially-decimated HMI magnetogram data often prevents the region from evolving before it can be correctly identified as a bipolar region. Using EUV only evolution, these cases typically become large enough to correctly identify in $<50$ iterations. Due to the presence of filament regions with the behavior seen in Figure 13b, we remove potential filament regions after evolving the contour for 50 EUV-only iterations. To allow for the variation in measured unipolarity seen in both examples in Figure 13 and in other similar cases, we remove regions with a unipolarity $U\geq 0.8$.
Interestingly, we additionally note in Figure 13 that the incorporation of magnetic field appears to result in faster convergence. While definitive reasons for this are unclear, we hypothesize that it is due to a combination of two factors. First, estimates of the CH are now more conservative in general since the evolution is guided away from QS regions by the unipolarity constraint. Second, there are now two forces (homogeneity and unipolarity) that are both acting on the contour, guiding it to a stable (albeit smaller) solution faster.
In summary, inclusion of HMI data and the unipolarity term in the energy functional constrains evolution of filament regions to remain small and isolated around the initial seed (which is determined based solely on EUV intensity). That constraint, however, can result in the area remaining too small for accurate statistics in the computation of unipolarity, resulting in misidentification of filaments as small CHs. These results indicate that a seed filtering process to remove filaments can utilize initial evolution with only EUV data, removal of regions with insufficient unipolarity, and evolution of remaining regions to convergence. This process is elaborated on in the next section.
@@ 5.1 Updated Pipeline
Based on the results from Section 4.2, which demonstrate that the initial seed does not provide sufficient information to correctly identify filaments, and the results from Section 3.4, which demonstrate that decimated HMI magnetogram data still constrains CH evolution to highly unipolar regions, we develop the following ACWE pipeline:
@@
    Mask off-disk areas and evolve using only the EUV observation(s) until the the unipolarity of each region can be accurately estimated, approximately 50 iterations. These initial iterations set $\lambda^{\rm{U}+}_{j}=\lambda^{\rm{U}-}_{j}=0$ where $j$ is the index of the HMI magnetogram data, and evolve using Equation 11 as the energy functional.
@@
    Continue evolving using both EUV and HMI magnetogram data (by setting $\lambda^{\rm{U}+}_{j}$ and $\lambda^{\rm{U}-}_{j}$ to non-zero values) until convergence.
We note that every ten iterations, both when evolving only EUV data and when evolving with both EUV and HMI magnetogram data, we reset the intensity of the off-disk regions to the channel-specific mean of the non-CH, on-disk region.
@@ 5.2 Results
For the results in this section, we use the EUV 193 Å observation, seed using $\alpha=0.3$, and perform EUV-only evolution for $50$ iterations using the homogeneity parameters from Section 3.3 before removing regions with a line-of-sight unipolarity metric $\geq 0.8$. We perform the full EUV+HMI magnetogram evolution on the remaining regions using the weights from Section 3.3 for all parameters. Figure 14 summarizes the similarity between the one-eighth-scale seeded and evolved EUV+HMI magnetogram segmentations in Section 3.4, and the segmentations generated here for CR 2133. We note that the median IOU and SSIM are both one ($1$) and that the median GCE and LCE are both zero (00), indicating that the vast majority of segmentations are identical. This is consistent with the expectation that seeding at one-eighth-scale already removed the majority of filaments (see Section 3.4). Figure 15 provides samples of the segmentations with differences between the methods. We note that the regions that are missing as a result of filtering the seed are regions that appear highly bipolar based on line-of-sight, full-scale magnetogram data. Referencing Figure 15c, we further note that this can, in rare cases, result in the loss of observations near disk edge. Exploration of a variable unipolarity threshold that is a function of distance from disk center is therefore an area of future research.
@@
Figure 14: Comparison between one-eighth-scale seeded and evolved EUV+HMI magnetogram segmentations and EUV+HMI magnetogram segmentations generated from a filtered seed over CR 2133. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
@@
Figure 15: Example of segmentations from CR 2133 generated from EUV and HMI magnetogram observations without filtering the initial seed (center) and with seed filtering (right). The caption under each example provides the unweighted line-of-sight unipolarity of the missing regions in order of right to left, top to bottom.
@@
Figure 16: Intersection Over Union between one-eighth-scale seeded and evolved EUV+HMI magnetogram segmentations, and EUV+HMI magnetogram segmentations generated from a filtered seed for each CR. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
(a) CR 2099
(b) CR 2100
(c) CR 2101
(d) CR 2133
Figure 17: Intersection Over Union between one-eighth-scale seeded and evolved EUV+HMI magnetogram segmentations, and EUV+HMI magnetogram segmentations generated from a filtered seed for each CR.
@@
Noting that IOU is the most stringent of the four metrics, we provide the comparisons between EUV+HMI magnetogram data with and without seed filtering for all for CRs as box plots in Figure 16 and as 200-bin histograms in Figure 17. Visual inspection of all segmentations with discrepancies revealed that the majority of missing regions consisted of small regions that appear highly bipolar in full-scale line-of-sight magnetogram data. The remaining missing regions consisted of regions along the disk edge. All regions that were retained, in every CR, appeared visually identical.

### v2

@@ 4.2 Determining Region Evolution
To determine if filaments can be removed during evolution, instead of as a post-hoc process, the unipolarity of every region as a function of number of iterations was calculated for both ACWE and QUACK segmentations. This was achieved by seeding and iterating at one-eighth-scale, saving a copy of the contour after each iteration, resizing the output to full-scale resolution, and calculating the LOS unipolarity of each region using the original, full-scale HMI magnetogram.
Figure 12 provides two examples of the change in measured unipolarity and change in area (measured in pixels at full-scale resolution) for a filament region. In the left panel of each example, we provide the LOS unipolarity, calculated via Equation 6, and a weighted unipolarity, weighting regions near disk center (where estimates of magnetic field strength are more accurate) using the method outlined in the SunPy documentation (The SunPy Community, 2024). In the right panels, we provide the area.
The examples in Figure 12 were chosen to demonstrate the two primary behaviors noted among filaments in CR 2133. The first case, seen in Figure 12a, consists of regions that quickly become bipolar as they evolve. We note that, when this occurs, the spatially-decimated magnetogram further constrains evolution as compared to an EUV-only evolution, preventing further growth of the region. In these cases the filament region can be identified and removed after very few iterations. This case is representative of most filaments in CR 2133.
@@
Figure 12: Region unipolarity (left) and region size in pixels (right) as a function of iteration for ACWE and QUACK segmentations generated at one-eighth scale resolution.
The second case, seen in Figure 12b, consists of regions that take several iterations before they can be correctly identified as filaments. In these cases the spatially-decimated magnetogram data often prevents the region from evolving before it can be correctly identified as bipolar. Using EUV-only evolution, these cases typically evolve enough to correctly identify in <50<50 iterations. Due to the presence of filaments with the behavior seen in Figure 12b, we remove potential filament regions after evolving the contour for 50 EUV-only iterations. To allow for the variation in measured unipolarity seen in both examples in Figure 12 and in other similar cases, we remove regions with a unipolarity $U\geq 0.8$.
Interestingly, we note in Figure 12 that the incorporation of magnetic field appears to result in faster convergence. While definitive reasons for this are unclear, we hypothesize that it is due to two factors. First, estimates of CHs are now more conservative since the contour is guided away from QS by the unipolarity constraint. Second, there are now two forces (homogeneity and unipolarity) acting on the contour, guiding it to a stable (albeit smaller) solution faster.
In summary, inclusion of HMI data and the unipolarity term in the energy functional constrains evolution of filament regions to remain small and isolated around the seed (which is determined based solely on EUV intensity). That constraint, however, can result in too small an area for accurate statistics in the computation of unipolarity, resulting in misidentification of filaments as small CHs. Thus, a seed filtering process to remove filaments can utilize initial evolution with only EUV data, removal of regions with insufficient unipolarity, and evolution of remaining regions to convergence using EUV and HMI data.
@@ 5.1 Updated Pipeline
Results in Section 4.2 demonstrate that the seed does not provide sufficient information to correctly identify filaments. Results in Section 3.4 demonstrate that decimated magnetograms still constrain CH evolution to highly unipolar regions. Based on these results, we develop the following QUACK pipeline:
@@
    Mask off-disk areas and evolve using only the EUV observation(s) until the the unipolarity of each region can be accurately estimated, approximately 50 iterations. These initial iterations set $\lambda^{\rm{U}+}_{j}=\lambda^{\rm{U}-}_{j}=0$ where $j$ is the index of the magnetogram data, and evolve using Equation 11 as the energy functional.
@@
    Continue evolving using both EUV and magnetogram data (by setting $\lambda^{\rm{U}+}_{j}$ and $\lambda^{\rm{U}-}_{j}$ to non-zero values) until convergence.
We note that every ten iterations, both when evolving only EUV data and when evolving with both EUV and magnetogram data, we reset the intensity of the off-disk regions to the channel-specific mean of the non-CH, on-disk region.
@@ 5.2 Results
In this section, unless otherwise stated, we use the EUV 193 Å observation, seed using $\alpha=0.3$, and perform EUV-only evolution for $50$ iterations using the homogeneity parameters from Section 3.3 before removing regions with a LOS unipolarity $U\geq 0.8$. We perform the full EUV+HMI magnetogram evolution on the remaining regions using the weights from Section 3.3.
5.2.1 Primary Results
@@
Figure 13: Comparison between one-eighth-scale QUACK segmentations with and without seed filtering for CR 2133. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
Figure 13 summarizes the similarity between the one-eighth-scale seeded and evolved QUACK segmentations with and without seed filtering for CR 2133. We note that the median IOU and SSIM are both $1$ and that the median GCE and LCE are both 0, indicating that the vast majority of segmentations are identical. This is consistent with the expectation that seeding at one-eighth-scale already removed the majority of filaments (see Section 3.4). Figure 14 provides samples of the segmentations with differences between inclusion and exclusion of seed filtering. We note that regions missing as a result of seed filtering are regions that appear highly bipolar based on LOS, full-scale magnetogram data. Referencing Figure 14c, we further note that this can, in rare cases, result in the loss of observations near disk edge. Exploration of a variable unipolarity threshold that is a function of distance from disk center is therefore an area of future research.
@@
Figure 14: Examples of QUACK segmentations from CR 2133 generated without filtering the initial seed (center) and with seed filtering (right). The caption under each example provides the unweighted LOS unipolarity of the missing regions in order of right to left, top to bottom.
Figure 15: IOU between one-eighth-scale QUACK segmentations with and without seed filtering for each CR. The box outlines the range between Q1 and Q3, with the median value in orange. The whiskers show 1.5 times the interquartile range. Outliers are marked with circles.
@@
Figure 16: IOU between one-eighth-scale QUACK segmentations with and without seed filtering for each CR.
@@
Noting that IOU is the most stringent of the four metrics, we provide comparisons between QUACK segmentations with and without seed filtering for all four CRs as box plots in Figure 15 and as 200-bin histograms in Figure 16. Both figures indicate that all segmentations with discrepancies appear to be outliers. Visual inspection of those segmentations with discrepancies reveal that the majority of regions missing when using seed filtering, but present when omitting seed filtering, consisted of small regions that appear highly bipolar in full-scale LOS magnetogram data. The remaining regions missing when using seed filtering consist of regions along the disk edge. All regions that were retained appear visually identical. We note that the plots in Figures 15 and 16 present data for four CRs; thus, while we analyze hundreds of segmentations per CR, the effective sample size is smaller in that CHs are expected to remain relatively stable in appearance over a single CR. We consider a larger sample size in Section 5.2.3.
5.2.2 Comparison to ACWE
In order to further compare QUACK to ACWE we utilize the dataset of Reiss et al. (2023). This dataset contains seven AIA Level 1 EUV images and the 45-second LOS HMI magnetograms with the closest observation time, all preprocessed to ensure alignment between the data products. The results in this section thus represent our first-ever segmentations using 45-second LOS HMI magnetograms (rather than 720-second magnetograms).
To provide a fair comparison between methods, segmentation is performed twice, once using a seed generated with $\alpha=0.3$, and once using $\alpha=0.4$, referred to as “QUACK03” and “QUACK04,” respectively. These two segmentations will be compared to “ACWE03” and “ACWE04” of Reiss et al. (2023), respectively, which begin with the exact same seed (with one exception). The two versions of ACWE in Reiss et al. (2023) were produced because a threshold of $\alpha=0.3$ does not yield any segmentation for the observation on 2 September 2016. Assuming that the full QUACK pipeline can mitigate erroneous detections (e.g., filaments) in subsequent steps, QUACK will increment $\alpha$ until a seed is generated. For this reason, a seed created with $\alpha=0.33$, automatically chosen by the algorithm, is used for this single observation in the QUACK03 segmentations.
(a) One filament removed, one filament reduced
(b) One filament reduced
(c) Example effect of minimizing growth into QS
Figure 17: Comparison between ACWE03 and QUACK03 segmentation for cases in the dataset of Reiss et al. (2023). In these examples, regions labeled as CH in Reiss et al. (2023) are blue and regions labeled as filaments in Reiss et al. (2023) are red.
Of the 5 filaments present in ACWE03, QUACK03 removed 3 filaments: one each from the observations on 6 June 2015, 11 August 2015, and 23 September 2015. The first retained filament, in the data from 23 September 2015, shown in Figure 17a, was reduced from $875$ pixels to $121$ pixels (at $512\times 512$ pixel resolution), equivalent to a reduction from $14293.863$ Mm2 to $2266.325$ Mm2. The second retained filament, in the data from 26 March 2016, shown in Figure 17b, was reduced from $591$ pixels to $51$ pixels ($512\times 512$ pixel resolution), a reduction from $8110.057$ Mm2 to $660.249$ Mm2. QUACK03 resulted in the loss of 9 additional CH regions, 8 at the limb and a faint CH near disk center on 4 May 2019. Of these cases, 8 CHs were removed due to seed filtering, including the case near disk center. Removal of CHs at disk edge is not surprising given that the LOS magnetic field less accurately portrays unipolarity at disk edge. In the final case, on 2 October 2014, shown in Figure 17c, the CH on the NW edge of the limb is absent from the seed and was identified in ACWE03 due to evolution through the QS from an adjacent CH. Since QUACK03 prevented growth into the QS, this CH is absent in QUACK03. This CH is also absent in ACWE04 and QUACK04.
Table 1: Comparison of filaments present in ACWE04 and QUACK04 segmentations for the datset of Reiss et al. (2023) showing the official label, number of pixels of each filament region at $512\times 512$ pixel resolution, and total area in Mm2.
Date Label ACWE04 Size QUACK04 Size
Pixels Mm2 Pixels Mm2
2014-07-15 Fil3 $523$ $7732.382$ $276$ $4049.278$
2015-01-21 Fil1 $214$ $6874.345$ $17$ $542.313$
2015-02-10 Fil1 $545$ $23922.795$ $95$ $6813.639$
2015-02-10 Fil4 $316$ $17609.477$ $26$ $814.206$
2015-02-10 Fil7 $1365$ $18021.584$ $671$ $8677.054$
2015-03-31 Fil2 $472$ $8455.888$ $43$ $652.559$
2015-04-18 Fil2 $185$ $6751.238$ $33$ $1400.311$
2015-08-11 Fil2 $1486$ $18654.977$ $297$ $3767.503$
2015-08-11 Fil3 $365$ $4972.121$ $48$ $641.910$
2015-09-23 Fil1 $2212$ $31573.298$ $176$ $2184.171$
2015-09-23 Fil2 $1712$ $25406.038$ $157$ $2255.122$
2015-09-23 Fil3 $300$ $4437.861$ $36$ $469.649$
2016-01-14 Fil2 $480$ $7811.111$ $27$ $411.945$
2016-03-26 Fil1 $687$ $9751.211$ $383$ $5216.005$
2016-03-26 Fil3 $197$ $3013.165$ $15$ $210.407$
2016-12-28 Fil2 $46$ $723.455$ $14$ $200.016$
2017-10-04 Fil2 $50$ $540.241$ 2121 $305.042$
2017-10-04 Fil3 $83$ $1967.337$ $32$ $747.750$
QUACK04 eliminated 6 of the 24 filaments found in ACWE04. Effects on the remaining filaments are summarized in Table 1. We note that our tally of filaments for ACWE04 differs from those in Reiss et al. (2023) due to a difference in our count of filaments on 26 March 2016. Seeding with $\alpha=0.4$ appears to have significantly reduced the effectiveness of seed filtering. Exploring the relationship between $\alpha$ and the unipolarity threshold used in seed filtering should thus be an area of future research. Seed filtering also resulted in the loss of 5 CH regions, all at disk edge. We provide a summary of ACWE and QUACK performance in Table 2. All areas are first-order approximations assuming the area of each pixel $A_{i}$ of the full-scale image is
$A_{i}=\frac{A_{\rm{nom}}}{\cos(\theta_{i})},$ (12)
where $A_{\rm{nom}}=0.189$Mm2 is the nominal area of an AIA pixel at disk center (Lemen et al., 2012) and $\theta_{i}$ is the angular distance from the Earth-Sun line.
Table 2: Comparison of methods for Reiss et al. (2023) dataset, outlining number of CHs correctly identified, number of CHs missed, number of filaments incorrectly segmented, number of filaments correctly ignored, and number of other object or features caught by each algorithm.
Method Coronal Holes Filaments Other
Identified Missed Identified Ignored Identified
ACWE03 71 15 5 66 2
QUACK03 63 23 2 69 2
ACWE04 74 12 24 47 12
QUACK04 69 18 18 53 10
5.2.3 Extended Study of Segmentation Characteristics
As additional validation, we use the dataset consisting of observations from the start of CR 2099 (13 July 2010) through the end of CR 2294 (2 March 2025), with a cadence of approximately 1 day between observations. We note, however, that a data gap of 8 days exists in CR 2180, a gap of 6 days in CR 2136, a gap of 4 days in CR 2207, and a gap of 2 days in CR 2243. In this extended dataset, segmentations are generated using the standard QUACK pipeline (Section 5).
Figure 18 summarizes the total area of each observation within the dataset, providing both the measured area of each segmentation in the form of a scatter plot and the mean area for sliding windows of various sizes (line plots). The measured area, provided in Mm2, is the first-order approximation estimated using Equation 12. We note that while the estimated CH area varies from one observation to the next, the means of the sliding windows indicate that CH area, as observed by our method, is lower during the two solar maxima and higher during the solar minimum. This is consistent with the expectation of fewer CHs during solar maximum. We further note that the current period of high solar activity is not present in either of the previous two datasets, and that no data after 22 February 2013 was used during the development of this algorithm, suggesting strong generalization of the QUACK pipeline.
Figure 18: Total area of all regions identified as CHs using QUACK for each observation in the daily cadence dataset (dots) and mean area of all segmentations in a sliding window centered on the specified date (lines).
For further study we divided each segmentation into the individual CH regions using the process outlined in Section 4.1. This ensures that any disjoint regions along a CH boundary can be correctly associated with their respective CH. We calculated the area of each identified region and the location of its center of mass. We also calculated the mean magnetic field density $\bar{B}$ by first correcting for projection effects in the LOS observation $B$ on a pixel-by-pixel basis using
$B_{i\mathrm{Corr}}=\frac{B_{i}}{\cos(\theta_{i})},$ (13)
then calculating the total unbalanced magnetic flux of each region using
$\Phi=\sum_{i}B_{i\mathrm{Corr}}A_{i},$ (14)
and finally defining mean magnetic field density as
$\bar{B}=\frac{\Phi}{\sum_{i}A_{i}}.$ (15)
For this, we utilized Level 1.5 HMI magnetograms, which have been aligned to the Level 1.5 AIA observations. This reprojection process was achieved by converting each magnetogram to a SunPy map object and using the reproject_to method (The SunPy Community et al., 2020). Alternatively, rescaling and aligning the segmentations to the Level 1 HMI magnetograms (and adjusting $A_{\rm{nom}}$ accordingly) produces highly similar results.
Figure 19: Total area of each region identified as a CH using QUACK for each observation in the daily cadence dataset (dots) and mean area of the individual regions in a sliding window centered on the specified date (lines). See Figure 18 for a legend.
Figure 20: Number of CHs found in each observation. Note that six dates exceed the range of [0,20] of this plot: 30 March 2014, 12 March 2015, 7 May 2019, 5 May 2024, 08 May 2024, and 09 May 2024. All of these cases, except 7 May 2019, have exposure times <0.15<0.15s.
(a) All data within $\pm 500^{\prime\prime}$
(b) Zoomed in, with sliding window means
Figure 21: Absolute mean magnetic field density ($|\bar{B}|$) of individual CHs identified using QUACK for each CH within $\pm 500^{\prime\prime}$ (dots) and in a sliding window centered on the specified date (lines). See Figure 18 for a legend of the line plots.
In Figure 19, we provide a plot of the area in Mm2 of all CH regions we identified. Excluding local peaks caused by disk-edge CHs, we note that the mean area (for each sliding window) is relatively consistent across the dataset. This strongly indicates that the increase in CH area during solar minimum seen in Figure 18 is predominantly driven by an increase in the total number of CHs. This is confirmed in Figure 20, which shows the total number of CH regions in each observation. Across the full dataset (including the cases at disk edge) we find that the mean area of individual CHs is $3.48\times 10^{4}$ Mm2 and the median area is $9.88\times 10^{3}$ Mm2. This is similar to, but slightly higher than, the average area of $2.39\times 10^{4}$ Mm2 reported in Hofmeister et al. (2017).
Figure 21 shows the absolute mean magnetic field density ($|\bar{B}|$) of the regions with a center of mass within $\pm 500^{\prime\prime}$ in the Helioprojective Cartesian system (dots) and means of sliding windows (lines). We note that the mean of $|\bar{B}|$, across all observations within the $\pm 500^{\prime\prime}$ region is $2.60\pm 3.04$ G. Compared to the values of $2.9\pm 1.9$G in Heinemann et al. (2019) and $2.97\pm 1.55$G in Hofmeister et al. (2017), we find a slightly lower average, indicating that QUACK is segmenting CHs with underlying magnetic field densities commensurate with expectations. A potential source of difference may be that QUACK does not rely on an intensity threshold for CH delineation as in Hofmeister et al. (2017); Heinemann et al. (2019). We further note a cyclical trend in $|\bar{B}|$ corresponding to the solar cycle (see Figure 21b), with slightly larger $|\bar{B}|$ around solar maximum and slightly lower $|\bar{B}|$ around solar minimum. This is consistent with observations in Hofmeister et al. (2017) (and references therein). The larger standard deviation found here as compared to Hofmeister et al. (2017); Heinemann et al. (2019) is most likely due to the few outliers seen in Figure 21a. These outliers are mainly concentrated in 2024 data, and correspond with significantly reduced exposure times, likely a result of solar flares associated with the Gannon storm. These outlier data are outside of the time range considered in Hofmeister et al. (2017); Heinemann et al. (2019) so it is impossible to state whether those intensity-based algorithms would be similarly affected. Eliminating May 2024 from the computation results in a mean of $|\bar{B}|$ of $2.50\pm 2.01$ G within $\pm 500^{\prime\prime}$.